OpenStudy (anonymous):

Rationalize the denominator of expression: sqrt (3) / (7)

OpenStudy (anonymous):

\[\sqrt{\frac{3}{7}}?\]

OpenStudy (anonymous):

The seven is correct, but it keeps telling me the three is wrong. So I'm sure what it's asking.

OpenStudy (anonymous):

if the question is \[\sqrt{\frac{3}{7}}\] then:\[\sqrt{\frac{3}{7}}=\frac{\sqrt{3}}{\sqrt{7}}=\frac{\sqrt{3}}{\sqrt{7}}*\frac{\sqrt{7}}{\sqrt{7}}=\frac{\sqrt{21}}{7}\]

OpenStudy (anonymous):

Okay. How do I enter the Square Root symbol online then? Because I'm doing homework online, and it keeps telling me the answer is incorrect.

OpenStudy (anonymous):

depends on the system u r using